How can 28,0000 be invested, part at 4% annual simple interest and the remainder at 10%, so that the interest earned by the two accounts is equal at the end of the year?
Answer by stanbon(75887)

How can 28,0000 be invested, part at 4% annual simple interest and the remainder at 10%, so that the interest earned by the two accounts is equal at the end of the year?
--------------------------
Let amt. invested at 4% be "x"; Interest is 0.04x dollars
---------------------------
Amt invsted at 10% is "28000-x" ; Interest is 0.1(28000-x)=2800-0.1x
----------------------------
EQUATION:
interest = interest
0.04x = 2800-0.10x
0.14x = 2800
x = $20000 (amount invested at 4%)
------------
28000-x = $8000 (amount invested at 10%)
=============================================
Cheers,
Stan H.